HUBBLE SPACE TELESCOPE DAILY REPORT #1756 PERIOD COVERED: 0000Z (UTC) 11/07/96 - 0000Z (UTC) 11/08/96 Daily Status Report as of 313/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 6188 (WF/PC-2 Cycle 5 Calibration: Dark) The WF/PC-2 was used to make internal dark calibration images. = The observations were completed as scheduled, and no problems were = reported. 1.2 Completed GHRS/2 6712 (The Gaseous Extent of Galaxies: How Far is = Too Far?) The High Resolution Spectrograph (Side Two) was used to make = spectrographic observations of the quasar Q0405-1219 located in a field = of galaxies in which proposers have carried out extensive redshift = surveys. A wavelength calibration image was also taken. The = observations were completed as planned and no problems were noted. 1.3 Completed WF/PC-2 6909 (WF/PC-2 Cycle 6 Calibration: Earth Flats) The WF/PC-2 was used to make observations of the bright Earth to = obtain streak flats which are used to construct high quality flat fields = for the WF/PC-2 filter set. The observations were executed as scheduled, = and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 3 Successful : 3 Scheduled Reacquisitions : 11 Successful : 11 2.2 FHST Updates: Scheduled : 8 Successful : 8 2.3 Operations Notes: There were no significant operational problems reported. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH